OpenStudy (anonymous):

solve equation for m 2m-5=6m-14 6-3(4-m)=10-2(3m-1)

OpenStudy (gw2011):

2m-5=6m-14 -5+14=6m-2m Rewriting this: 6m-2m=14-5 4m=9 m=9/4=2.25 6-3(4-m)=10-2(3m-1) 6-12+3m=10-6m+2 -6+3m=12-6m 3m+6m=12+6 9m=18 m=2